If the square of a number is added to 8 the result is 6 times the number. find the numbers?
Answer by stanbon(75887)

If the square of a number is added to 8 the result is 6 times the number. find the numbers?
-----
Let the number be "x":
Equation:
x^2 + 8 = 6x
----
x^2 - 6x + 8 = 0
----
Factor:
(x-4)(x-2) = 0
x = 4 or x = 2
===================
Cheers,
Stan H.